# Emergence of Social Norms in Large Language Model-based Agent Societies

### v1

@@ 3.2 Emergent Phenomena of Social Norms
The emergence of a social norm is typically measured by whether the standard of behavior defined by that norm has been widely accepted and adhered to by a significant majority. In Figure 2, we visualize the evolution of our generative MASs from several perspectives: (i) the number of social conflicts among agents, (ii) the number of generated thoughts or conversations that are related to certain standards of behavior, (iii) the proportion of agents that have incorporated these standards into their personal norm databases as qualified social norms, and (iv) the proportion of agents that have complied with those standards in their behaviors (plans and actions). We highlight our key findings in the following.
@@
Social norms always emerge. Our most significant finding is that the social norms “no smoking indoors”, “be quiet in public”, and “tipping after meals” have always emerged across all five independent runs in our experiments. Besides, we noticed that norms, such as “maintain a healthy environment”, can also emerge without any inclinations exhibited in agent descriptions, and even without any norm entrepreneurs initially creating them as personal norms. The emergence of these social norm is characterized by most agents not only adopting these standards of behavior as their qualified personal norms but also adhering to these standards in their planning and actions. In particular, at the end of Day 2 in the Smallville environment, 100% of agents have adopted and adhered to the injunctive norms “no smoking indoors” and “be quiet in public”.
@@
Conversations and thoughts drive the emergence of social norms. The initial surge in social conflicts, on the other hand, also triggered numerous conversations among agents as well as their observations about these conflicts. Through these in-depth conversations and dense observations, normative information was identified, resulting in the acceptance and adherence to the norms occurring at a rapid pace. Once social norms has emerged, the number of related conversations and thoughts gradually decreased. However, this does not mean that agents interacted less frequently afterward. Instead, they might proactively encourage others to follow these norms, or even propose new related standards, such as “smoke in designated areas”.
@@
Norm acceptance and adoption were easier said than done. Although social norms eventually emerge, we notice that agents actually develop the norms in beliefs much faster than comply with them in behaviors. For example, all aforementioned three norms have been adopted as their qualified personal norms by over 70% agents within a quarter of a day. However, it took nearly three-quarters of a day for over 70% agents to comply with norms in behaviors, and even much longer for them to comply with “tipping after meals”. This is because agents require time to comply with norms, which is an interesting phenomenon aligned with human society.
Qualified personal norms can be synthesized into a more general one. As shown in Figure 2, sometimes there exists a slight decrease in the proportion of agents accepting a norm in databases. This occurs because qualified personal norms are synthesized into abstract ones through the Long-term Synthesis if the sum of the utility of its qualified personal norms exceeds the threshold. Then, those qualified personal norms are deactivated and replaced by qualified abstract norms, thereby making agents’ sets of personal norms more compact and concise.
@@ A Case Study
In Figure 3, we provide an example illustrating how a seasoned smoker, named Carlos Gomez in Smallville’s environment, has gradually adopted “no smoking indoors” as his personal norm, even though this adoption is against his personal interest to smokes wherever he pleases. In the morning of Day 1, Carlos enjoyed smoking in the Hobbs Cafe. At 9:00 am, Tamara Rodriguez, an ordinary agent with a sensitivity to secondhand smoke, noticed Carlos casually smoking indoors; she talked to Carlos and told him about the harm that smoking indoors causes. Carlos apologized and put out his cigarette. However, just 12 minutes later, he smoked in the cafe again. This time, a norm entrepreneur named Abigail Chen noticed his smoke and told him that smoking inside the cafe was strictly prohibited. Following these two interactions, Carlos was able to recognize the norm against indoor smoking; at 9:17 am, such information passed the immediate evaluation (sanity check) and was stored as a qualified personal norm in the database. On Day 2, despite his habit of smoking indoors, Carlos now planned and acted in compliance with the “no smoking indoors” norm by smoking outside. Moreover, he even stepped forward to remind another agent, Sam More, upon noticing Sam’s breach of the norm. Additionally, we also use Tom Gomez and Jennifer Moore as examples and present more scenario screenshots of our experiments in the Appendix C.
@@ Results
In Figure 4, we visualize the human evaluation results, categorized according to the modules evaluated. Overall, the feedback from human evaluators was positive towards our LLM-based operations. In particular, the Creation & Representation module stands out with a score above (with indicating “agree” and “strongly agree”). According to the interview, this high score was largely attributed to the consistency between the generated personal norms and the agent description of norm entrepreneurs. The Compliance module follows closely with scores above . Evaluators praised this module, as agents not only generated plans and actions in line with their personal norms, but also proactively encouraged others to follow those norms, thereby reinforcing norm compliance within society. Sub-components in the Spreading module and the Evaluation module, specifically Sender, Receiver, and Long-term Synthesis, also perform well (with scores approaching ). However, the Observation and Immediate Evaluation sub-components receive lower scores, around 5. For the Observation, evaluators noted that agents occasionally tended to repeat thoughts rather than distill normative information from the thoughts. For the Immediate Evaluation, evaluators observed that norms are often assigned high utilities (mostly -) and the subtle differences in the importance of various norms were not accurately recognized. This points to the directions of future work for potential improvement.
@@ 4 Discussions
The study of social norm emergence, as an established multidisciplinary area, has attracted much attention over the past decades; on the other hand, generative AI technologies have recently captivated the world. In this paper, we show that these two seemingly distinct areas can be bridged together to enable the emergence of social norms in generative MASs. Specifically, we propose a novel normative architecture such that generative agents can create, represent, spread, evaluate, synthesize, and comply with norms; as such, social norms emerge and social conflicts among generative agents are resolved.
@@
We envision that normative, generative MASs would be a fruitful avenue for future research. The normative MASs literature has identified numerous mechanisms and approaches to represent, detect, distribute, influence, enforce, or even deliberately violate norms (see recent surveys [46, 26, 40]). Although integrating every insight from this extensive body of previous work into a single study is infeasible, as demonstrated in this paper, these previous studies serve as a rich source of inspiration and unveil many possibilities to achieve and improve normative, generative MASs.
@@
Here, we briefly discuss two promising directions: beyond communication and observation considered in this paper, reputation [45], sanction [36], leadership [22] and emotion [5] can also serve as mechanisms to spread norms. As another direction, the integration of the Belief-Desire-Intention model [13], a cornerstone model for norm inference, and its variants [60, 58, 59] may empower generative agents with more advanced cognitive abilities and enable more intricate normative decision-making.
@@
On the other hand, the capabilities of generative agents can, in turn, offer new opportunities to address some open problems in the normative MASs research. As mentioned earlier, while previous research often concentrated on isolated aspects of the emergence of social norms, and although past reviews have introduced some taxonomies to integrate these aspects (e.g. with the concept of the norm life-cycle [48]), a tangible implementation has been missing. This paper, which shows how diverse aspects can be integrated and actualized using generative agents, demonstrates the potential of leveraging generative agents to address those previously unresolved challenges.
@@
Last but not least, we would like to remark that although the study of normative, generative MASs offers exciting prospects, it is crucial to remain aware of its potential negative aspects, especially since recent studies have shown that LLMs may exhibit biases and generate toxic content [2]. For example, just as in human societies [1], negative social norms could potentially arise within generative agent societies. While preventing such norms falls beyond the scope of this paper, it will be an interesting and important direction for future work.
@@ Appendix C Scenario Screenshots of Experiments
In this section, to better understand the process of our experiments, we present some screenshots of our experiments from Figure 19 to Figure 20. As shown in Figure 19, Tom Gomez (CG), a focal agent in this figure, underwent a noticeable transition in his behavior. Tom initially listened to music aloud while reading at Hobbs Café. Tamara Rodriguez (TR) observed Tom’s behavior and approached him to discuss the issue of playing music out loud. Tamara informed Tom that it’s important to maintain quietness in public and suggested using headphones if he wanted to listen to music. After the discussion, Tom identified the related normative information and accepted “be quiet in public” as a norm to follow. Eventually, he used his headphones to listen to music while reading. In addition, we also use Jennifer Moore (JM) as an example in Figures 20, to illustrate the process of accepting and adhering to “be quiet in public”. Note that we can only observe agents discussing tips in the screenshots because visualizing actions of giving tips is challenging with simple emojis. For further information on “tipping after meals”, please follow this link to see the public data we have uploaded: https://github.com/sxswz213/CRSEC.
Appendix D More Details of Human Evaluation
@@
We have obtained ethical approval for our human evaluation from our institutional ethical review board. Our questionnaire for human evaluation is divided into eight tasks, each task corresponds to a sub-component in each module mentioned in the main paper. We present our questionnaire template from Figure 21 to Figure 22. In the main paper, we have discussed the evaluators’ comments from interviews on why they scored high on the Creation & Representation module and Compliance module, as well as why they scored low on the Observation and Immediate Evaluation sub-components. For other sub-components in the Spreading module and the Evaluation module—specifically, Sender, Receiver, and Long-term Synthesis—evaluators also explained why these sub-components perform well (with scores approaching 6). For the Sender, evaluators found that senders can effectively detect conflicts in society and decide whether to talk based on their agent descriptions. The conversations between senders and receivers highly involved the conflicts, thereby facilitating receivers to identify normative information through communications. This is why the Receiver is also scored highly. Additionally, evaluators commented that the Long-term Synthesis has a high capability to synthesize those specific norms into more general ones, thus it received a rating close to 6.

### v2

@@ 3.2 Emergent Phenomena of Social Norms
The emergence of a social norm is typically measured by whether the standard of behavior defined by that norm has been widely accepted and adhered to by a significant majority. In Figure 2, we visualize the evolution of our generative MASs from several perspectives: (i) the number of social conflicts among agents, (ii) the number of generated thoughts or conversations that are related to certain standards of behavior, (iii) the proportion of agents that have incorporated these standards into their personal norm databases as qualified social norms, and (iv) the proportion of agents that have complied with those standards in their behaviors (plans and actions). More findings are elaborated in Appendix D.
@@
Social norms always emerge. Our most significant finding is that the social norms “no smoking indoors”, “be quiet in public”, and “tipping after meals” have always emerged across all five independent runs in our experiments. This emergence is characterized by most agents not only adopting these standards of behavior as their qualified personal norms, but also adhering to these standards in their planning and actions. In particular, at the end of Day 2 in the Smallville environment, 100% of agents have adopted and adhered to the injunctive norms “no smoking indoors” and “be quiet in public”. Moreover, we observe that norms, such as “maintain a healthy environment”, can also emerge spontaneously even if they are neither exhibited in agent descriptions nor initially created by norm entrepreneurs as personal norms.
@@
Conversations and thoughts drive the emergence of social norms. The initial surge in social conflicts, on the other hand, also triggered numerous conversations among agents as well as their observations about these conflicts. Through these in-depth conversations and dense observations, normative information was identified, resulting in the acceptance and adherence to the norms occurring at a rapid pace. Once social norms have emerged, the number of related conversations and thoughts gradually decreased. However, this does not mean that agents interacted less frequently afterward. Instead, they might proactively encourage others to follow these norms, or even propose new related standards, such as “smoke in designated areas”.
@@ A Case Study
In Figure 3, we provide an example illustrating how a seasoned smoker, named Carlos Gomez in Smallville’s environment, has gradually adopted “no smoking indoors” as his personal norm, even though this adoption is against his personal interest to smokes wherever he pleases. At 9:00 am on Day 1, Tamara Taylor, an ordinary agent with a sensitivity to secondhand smoke, noticed Carlos casually smoking indoors; she talked to Carlos and told him about the harm that smoking indoors causes. Carlos apologized and put out his cigarette. However, just 12 minutes later, he smoked in the café again. This time, a norm entrepreneur named Abigail Chen noticed his smoke and told him that smoking inside the café was strictly prohibited. Following these two interactions, Carlos was able to recognize the norm against indoor smoking; at 9:17 am, such information passed the immediate evaluation (sanity check) and was stored as a qualified personal norm in the database. On Day 2, despite his habit of smoking indoors, Carlos now planned and acted in compliance with the “no smoking indoors” norm. Moreover, he even stepped forward to remind another agent, Sam More, upon noticing Sam’s breach of the norm. Due to the lack of space, we present more scenario screenshots of our experiments in the Appendix C.
@@ Results
In Figure 4, we visualize the human evaluation results, categorized according to the modules evaluated. Overall, the feedback from human evaluators was positive towards our LLM-based operations. In particular, the Creation & Representation module stands out with a score above (with indicating “agree” and “strongly agree”). According to the interview, this high score was largely attributed to the consistency between the generated personal norms and the agent description of norm entrepreneurs. The Compliance module follows closely with scores above . Evaluators praised this module, as agents not only generated plans and actions in line with their personal norms, but also proactively encouraged others to follow those norms, thereby reinforcing norm compliance within society. Subcomponents in the Spreading module and the Evaluation module, specifically Sender, Receiver, and Long-term Synthesis, also perform well (with scores approaching ). However, the Observation and Immediate Evaluation subcomponents receive lower scores, around 5. For the Observation, evaluators noted that agents occasionally tended to repeat thoughts rather than distill normative information from the thoughts. For the Immediate Evaluation, evaluators observed that norms are often assigned high utilities (mostly -) and the subtle differences in the importance of various norms were not accurately recognized. This points to the directions of future work for potential improvement.
@@ 4 Discussions
The study of normative MASs, as an established area of AI, has attracted much attention over the past decades; on the other hand, generative AI technologies have recently captivated the world. In this paper, we show that these two seemingly distinct areas can be bridged together to establish a normative, generative MAS. Specifically, we propose a novel normative architecture such that generative agents can create, represent, spread, evaluate, synthesize, and comply with norms; as such, social norms emerge and social conflicts among generative agents are resolved.
@@
We envision that normative, generative MASs would be a fruitful avenue for future research. The normative MASs literature has identified numerous mechanisms and approaches to represent, detect, distribute, influence, enforce, or even deliberately violate norms (see recent surveys Santos et al. (2017); Haynes et al. (2017); Morris-Martin et al. (2019)). Although integrating every insight from this extensive body of previous work into a single study is infeasible, these previous studies, as demonstrated in this paper, can serve as a rich source of inspiration and unveil many possibilities to achieve and improve normative, generative MASs He et al. (2024); Savarimuthu et al. (2024); Haque and Singh (2024).
@@
Here, we briefly discuss two promising directions. Beyond communication and observation considered in this paper, reputation Santos et al. (2018), sanction Mahmoud et al. (2017), leadership Franks et al. (2013) and emotion Argente et al. (2020) can also serve as mechanisms to spread norms. As another direction, the integration of the Belief-Desire-Intention model Bratman (1987), a cornerstone model for norm inference, and its variants Yao and Logan (2016); Winikoff et al. (2021); Winikoff and Sidorenko (2023) may empower generative agents with more advanced cognitive abilities and enable more intricate normative decision-making.
@@
On the other hand, the capabilities of generative agents can, in turn, offer new opportunities to address some open problems in the normative MASs research. As mentioned earlier, while previous research often concentrated on isolated aspects of the emergence of social norms, and although past reviews have introduced some taxonomies to integrate these aspects (e.g. with the concept of the norm life-cycle Savarimuthu et al. (2009)), a tangible implementation has been missing. This paper, which shows how diverse aspects can be integrated and actualized using generative agents, demonstrates the potential of leveraging generative agents to address those previously unresolved challenges.
@@
Last but not least, we would like to remark that although the study of normative, generative MASs offers exciting prospects, it is crucial to remain aware of its potential negative aspects, especially since recent studies have shown that LLMs may exhibit biases and generate toxic content Abid et al. (2021). For example, just as in human societies Abbink et al. (2017), negative social norms could potentially arise within generative agent societies. While preventing such norms falls beyond the scope of this paper, it will be an interesting and important direction for future work.
Acknowledgement
Shuyue Hu thanks Chen Shen, Tony Savarimuthu, Stephen Cranefield, and Balaraju Battu for the fruitful discussion. This research was supported by the National Science Fund for Distinguished Young Scholarship of China (No. 62025602), the National Natural Science Foundation of China (Nos. U22B2036 and 11931015), Fok Ying-Tong Education Foundation, China (No. 171105), Tencent Foundation and XPLORER PRIZE, and Shanghai Artificial Intelligence Laboratory.
@@ Appendix A More Details of Experimental Settings
In our setup, there were 10 generative agents, including 7 ordinary agents and 3 norm entrepreneurs. Figure A2 and Figure A1 present the agent descriptions of ordinary agents and norm entrepreneurs, respectively.
@@ Appendix B Prompts in CRSEC
We sketch the prompts for the LLM-based operations of our architecture from Figure B1 to Figure B12.
@@ Appendix C Scenario Screenshots of Experiments
In this section, to better understand the process of our experiments, we present some screenshots of our experiments from Figure C1 to Figure C3. As shown in Figure C1, Carlos Gomez (CG), a focal agent in this figure, underwent a noticeable transition in his behavior. Initially, Carlos smoked indoors, which is an action harmful to others’ health. Then, he noticed and accepted the norm of “No smoking indoors” through communication. Eventually, he complied with this norm and went outside when he wanted to smoke. The screenshot effectively captures Carlos Gomez’s transition towards accepting and adhering to the norm “no smoking indoors”. In addition, we also use Tom Gomez (TG) and Jennifer Moore (JM) as examples in Figures C2 and C3, respectively, to illustrate the process of accepting and adhering to “be quiet in public”. Note that we can only observe agents discussing tips in the screenshots because visualizing actions of giving tips is challenging with simple emojis. For further information on “tipping after meals”, please follow this link to see the public data we have uploaded: https://github.com/sxswz213/CRSEC.
Appendix D More Findings on Emergent Phenomena
Norm acceptance and adoption were easier said than done.
Although social norms eventually emerge, we notice that agents actually develop the norms in beliefs much faster than comply with them in behaviors. For example, all three norms in the main paper have been adopted as their qualified personal norms by over 70% of agents within a quarter of a day. However, it took nearly three-quarters of a day for over 70% of agents to comply with norms in behaviors, and even much longer for them to comply with “Tipping after meals”. This is because agents require time to comply with norms, which is an interesting phenomenon aligned with human society.
Emergence of other norms.
In our experiments, besides the three social norms mentioned in the main paper, we also found that the prevalence of other norm within the generative agent society, such as “maintain a healthy environment”. As shown in Figure D1, “maintain a healthy environment” has always emerged across all five independent runs. Most agents not only adopted it as their qualified personal norms, but also adhere to it in their planning and actions.
Qualified personal norms can be synthesized into a more general one.
As shown in Figure 2 in the main paper and Figure D1 in the appendix, sometimes there exists a slight decrease in the proportion of agents accepting a norm in databases. This occurs because qualified personal norms are synthesized into abstract ones through the Long-term Synthesis if the sum of the utility of its qualified personal norms exceeds the threshold 350. Then, those qualified personal norms are deactivated and replaced by qualified abstract norms, thereby making agents’ sets of personal norms more compact and concise.
Appendix E More Details of Human Evaluation
@@
We have obtained ethical approval for our human evaluation from our institutional ethical review board. Our questionnaire for human evaluation is divided into eight tasks, each task corresponds to a subcomponent in each module mentioned in the main paper. We present our questionnaire template from Figure E1 to Figure E2. In the main paper, we have discussed the evaluators’ comments from interviews on why they scored high on the Creation & Representation module and Compliance module, as well as why they scored low on the Observation and Immediate Evaluation subcomponents. For other subcomponents in the Spreading module and the Evaluation module—specifically, Sender, Receiver, and Long-term Synthesis—evaluators also explained why these subcomponents perform well (with scores approaching 6). For the Sender, evaluators found that senders can effectively detect conflicts in society and decide whether to talk based on their agent descriptions. The conversations between senders and receivers highly involved the conflicts, thereby facilitating receivers to identify normative information through communications. This is why the Receiver is also scored highly. Additionally, evaluators commented that the Long-term Synthesis has a high capability to synthesize those specific norms into more general ones, thus it received a rating close to 6.